# Injective envelopes of $C^{*}$-algebras as maximal rigid multiplier covers

### v2

@@ 4. Main theorem: $M(E_{\max})$ realises $I(A)$
Set $B:=M(E_{\max})$. By Definition 3.1, the inclusion $A\hookrightarrow B$ is rigid. By Lemma 3.10, it is essential in the quotient sense. By Theorem 2.1, $B$ is (canonically) the injective envelope of AA; that is, $B\cong I(A)$ over $A$. ∎
@@
Hamana showed that $I(A)$ need not be a von Neumann algebra; it is in general only monotone complete [7]. Our framework is compatible with this: nothing in the definition of a multiplier cover forces $M(E)$ to be $W^{*}$, and, as noted in Remark 2.3, even if one works inside a $W^{*}$ ambient algebra $B$, one typically gets a quotient $B/\mathrm{Ann}(J)$, which need not be $W^{*}$.
@@ 5. The commutative case revisited
    First, we the use Kuratowski–Zorn lemma to get a maximal regular refinement $T^{\ast}$ of the topology of $X$ for which $\mathrm{id}\colon(X,T^{\ast})\to(X,T)$ is irreducible. The resulting $(X,T^{\ast})$ is completely regular and extremally disconnected [1].
@@
To realise $I(C(X))$ as a multiplier algebra, work after compactification. Choose any dense open set $U\subseteq G(X)$ (for instance, an increasing union of clopen sets). Then
$M\big(C_{0}(U)\big)\ \cong\ C_{b}(U)\ \cong\ C(\beta U)\ =\ C(G(X)),$
since $\beta U$ identifies with the closure of $U$ inside $G(X)$, which equals $G(X)$ by density. Thus
@@ 6. Examples and remarks
If one works in a unital ambient algebra $B$ and chooses $J\triangleleft B$, then $M(J)\cong B/\mathrm{Ann}(J)$ (Remark 2.3). Thus $M(J)$ equals $B$ only when $J$ is essential in $B$. In particular, taking $B=\prod_{i}\mathcal{B}(H_{i})$ does not force $M(J)$ to be all of $B$ unless $J$ is essential; and even quotients of $W^{*}$-algebras need not be $W^{*}$ (for example the Calkin algebra is not). This removes the apparent tension with the fact that $I(A)$ is not generally von Neumann [7].
Acknowledgements
We are indebted to Aleksander Błaszczyk for his elegant topological construction [1], which inspired the present point of view.

### v3

@@ 4. Main theorem: $M(E_{\max})$ realises $I(A)$
Set $B:=M(E_{\max})$ and let $j:A\hookrightarrow I(A)$ be the Hamana embedding.
By injectivity of $I(A)$ as an operator system, the inclusion $j:A\to I(A)$ extends to a ucp map
$\Theta:B\longrightarrow I(A)\qquad\text{with}\qquad\Theta|_{A}=j.$
Thus $(E_{\max},\iota_{\max})\preccurlyeq(I(A),j)$ (viewing $(I(A),j)$ as a rigid $A$-multiplier cover).
Since $(E_{\max},\iota_{\max})$ is maximal among rigid covers, we also have $(I(A),j)\preccurlyeq(E_{\max},\iota_{\max})$. Hence there exists an $A$-preserving ucp map
$\Lambda:I(A)\longrightarrow B\qquad\text{with}\qquad\Lambda\circ j=\iota_{\max}.$
Then $(\Lambda\circ\Theta)|_{A}=\mathrm{id}_{A}$, so by rigidity of the cover $(E_{\max},\iota_{\max})$ we obtain
$\Lambda\circ\Theta=\mathrm{id}_{B}.$
In particular, $\Theta$ is a unital complete order embedding of $B$ into $I(A)$ which is the identity on $A$.
Moreover, $\Theta\circ\Lambda:I(A)\to I(A)$ is a ucp idempotent fixing $A$ pointwise, hence $\Theta(I(A))=\Theta(B)$ is an injective operator subsystem of $I(A)$ containing $j(A)$. By the minimality property in Theorem 2.1(a), we must have $\Theta(B)=I(A)$. Therefore $\Theta$ is surjective, and hence a complete order isomorphism over $A$.
Since $B$ and $I(A)$ are $C^{\ast}$-algebras, a unital complete order isomorphism is automatically a *-isomorphism. Thus $M(E_{\max})=B\cong I(A)$ canonically over $A$. ∎
@@
Hamana showed that $I(A)$ need not be a von Neumann algebra; it is in general only monotone complete [7]. Our framework is compatible with this: nothing in the definition of a multiplier cover forces $M(E)$ to be $W^{*}$, and, as noted in Remark 2.3, even if one works inside a $W^{*}$ ambient algebra $B$, the canonical map $B\to M(J)$ need not be surjective.
@@ 5. The commutative case revisited
    First, we use the Kuratowski–Zorn lemma to get a maximal regular refinement $T^{\ast}$ of the topology of $X$ for which $\mathrm{id}\colon(X,T^{\ast})\to(X,T)$ is irreducible. The resulting $(X,T^{\ast})$ is completely regular and extremally disconnected [1].
@@
To realise $I(C(X))$ as a multiplier algebra, work after compactification. Choose a dense cozero set $U\subseteq G(X)$ (equivalently, a dense open $F_{\sigma}$ set; for instance, an increasing union of clopen sets). Then $U$ is locally compact (being open in the compact space $G(X)$), so
$M\big(C_{0}(U)\big)\ \cong\ C_{b}(U)\ \cong\ C(\beta U).$
Since $G(X)$ is extremally disconnected compact, it is an $F$-space, and cozero sets are $C^{\ast}$-embedded. Consequently every bounded continuous function on $U$ extends uniquely to $G(X)$, so restriction induces an isomorphism $C(G(X))\cong C_{b}(U)$. Equivalently, $\beta U\cong G(X)$ (because $U$ is dense and $C^{\ast}$-embedded). Thus
$M\big(C_{0}(U)\big)\ \cong\ C(G(X)).$
Thus
@@ 6. Examples and remarks
If one works in a unital ambient algebra $B$ and chooses $J\triangleleft B$, the canonical map $B\to M(J)$ has kernel $\mathrm{Ann}(J)$ and is injective iff $J$ is essential, but is seldom surjective (for instance, $J=\mathcal{K}(H)$ inside its unitisation). Thus even when $B$ is $W^{*}$, $M(J)$ need not be $W^{*}$. This removes the apparent tension with the fact that $I(A)$ is not generally von Neumann [7].